OpenStudy (anonymous):

Given the function f described by f(x) = x + 14, find each of the following: f(0) = ? f(-19) = ? f(-14) = ? f(16) = ? and f(d+1) = ?

OpenStudy (anonymous):

I think the first one would be 14, right?

OpenStudy (anonymous):

second one would be -5??

OpenStudy (anonymous):

Third one would be 0??

OpenStudy (anonymous):

next one would be 30?

OpenStudy (anonymous):

last one I have no clue...LOL

OpenStudy (aroub):

f(x)=x+14 f(0)=0+14=14 f(-19)=-19+14=-5 f(-14)=-14+14=0 f(16)=16+14=30 f(d+1) = d+1+14= =d=-14-1 =d=-15

OpenStudy (anonymous):

so last one would be -15?

OpenStudy (aroub):

all answers u got are true!! and the last one i think it will be -15

OpenStudy (anonymous):

WHOO HOOOOO!!!!

OpenStudy (aroub):

LOL! :)

OpenStudy (aroub):

thanks for the medal!! :)

OpenStudy (anonymous):

No problem...you deserve it. You have helped me TONS of times.

OpenStudy (anonymous):

BUT....I do have one you could help me with that I asked earlier and didn't get an answer. Its the one about the athletes salaries...

OpenStudy (aroub):

oh wait ill check but i might not help in this sorryy! but i will sure take a look at it!! :)

OpenStudy (anonymous):

thank you...

OpenStudy (aroub):

okay....

OpenStudy (aroub):

for the first one i have no idea!! sorrry! :( for the second and the third i think u got the answers! / and if u want repost the question again!! im soorryyyy!! :(

OpenStudy (anonymous):

okay, I appreciate your help more than you know!! I will repost it.